OpenStudy (anonymous):

minimum or maximum of 2x^2+4x-18

OpenStudy (anonymous):

it is a parabola that opens up (because the leading coefficient 2 is positive) therefore it has a minimum, not a maximum 

OpenStudy (anonymous):

the minimum value is the second coordinate of the vertex the first coordinate of the vertex of \[y=ax^2+bx+c\] is \[-\frac{b}{2a}\] you have \[y= 2x^2+4x-18\] so the first coordinate of the vertex is \[-\frac{4}{2\times 2}=-1\]

OpenStudy (anonymous):

the second coordinate of the vertex is what you get when you replace \(x\) by \(-1\)

OpenStudy (dan815):

\[y=a(x^2+bx/a ) + c\\ y=a(x+b/2a)^2+c-a(b/2a)^~completing ~square\\ y=f(x+b/2a) form, if~base ~is ~y=x^2,~then ~this~means~shift~b/2a~to ~left\\ x=-b/2a,x-coord~point~of~vertex\\ y=a(x+b/2a)^2 +c-a(b/2a)^2\] y=f(x-k)+B form, y=x^2+b, we know vertex at 0,b similarily for y=f(x-k)+B, vertex at k,b y=a(x+b/2a)^2 +c-a(b/2a)^2, vertex at (-b/2a,c-a(b/2a)^2)

OpenStudy (dan815):